OpenStudy (anonymous):

Write y = –1/5x – 3 in standard form using integers.

OpenStudy (aravindg):

standard form will of a line will be like \[Ax+By+c=0\]

OpenStudy (aravindg):

so first step is to multiply both sides of equation by 5

OpenStudy (aravindg):

tell me what you get ?

OpenStudy (anonymous):

y = -5x + 15 ?

jimthompson5910 (jim_thompson5910):

multiply every term by the LCD 5 multiply y by 5 to get 5y multiply -1/5x by 5 to get -x multiply -3 by 5 to get -15 ------------------------------------- so after all of those multiplications, we go from y = –1/5x – 3 to 5y = -x - 15 Now get it into Ax + By = C form